OpenStudy (anonymous):

Calculate the speed of an electron that is accelerated from rest through a potential difference of 271V. Answer in units of m/s.

OpenStudy (mikurout):

9756643.716 m/s

OpenStudy (anonymous):

The electron is at rest so initially kinetic energy of e = 0 let the velocity of e is v then final kinetic energy of e = 1/2 * m * v^2 Change in kinetic energy = 1/2*m*v^2 But this kinetic energy is imparted to the electron due to potential difference. So if the Potential diff is V then imparted energy to e is e*V So, 1/2*m*v^2 = e*V Put e= 1.6*10^-19, m = 9.1*10^-31 and solve for v. Hope this helps.

OpenStudy (anonymous):

velocity=underoot 2ev/m,m=mass